OpenStudy (anonymous):

The formula d=rt relates the distance(d), the rate(r), and the time(t). Solve this equation for t, and use the rewritten equation to find an expression for the time it takes to travel a distance of (b^3-5b^2+7b+13) miles at a rate of (b+1) mph. It takes ?? hours to travel the distance?

OpenStudy (anonymous):

d=rt divide both sides by r t = d/r t = (b^3-5b^2+7b+13) / (b+1) now do the long division

OpenStudy (anonymous):

can you do long division of polynomials?

OpenStudy (anonymous):

I am trying but basically just winging it

OpenStudy (anonymous):

b^2 -6b + 13 ------------------------- b+1) b^3-5b^2+7b+13 b^3+b^2 -------- -6b^2+7b -6b^2 -6b --------- 13b + 13 13b + 13 --------- the result is b^2 -6b + 13 - this is the time taken in terms of b

OpenStudy (anonymous):

wow, you are good. thnx!